OpenStudy (mallorysipp234):

What is the amplitude, period, and phase shift of f(x) = −4 sin(2x + π) − 5?

OpenStudy (mathmale):

Compare your f(x) = −4 sin(2x + π) − 5 to: y = a*sin (bx + c) + d Your a value is -4; your b is ? your c is ? your d is ? the amplitude of your sine wave is |a|, the absolute value of a. The period is 2Pi/b what is the horizonal phase shift? What is the vertical offset?

OpenStudy (sidsiddhartha):

use this y=Asin(Bx-C),A=amplitude, Period=2pi/B, Phase shift=C/B

OpenStudy (sidsiddhartha):

thanks but mathmale deserves the medal :)

OpenStudy (mallorysipp234):

so amplitude is -4, Period is pi im confused on phase shift @sidsiddhartha @mathmale

OpenStudy (mallorysipp234):

is phase shift pi/2

OpenStudy (sidsiddhartha):

yes i think so

OpenStudy (mathmale):

Actually, the phase shift is -c/b. But you're close!

OpenStudy (mathmale):

There's also a vertical offset. Your value of d is -5. that means you graph the sine function first, and then, when you're done, you shift the whole graph downward 5 units.

OpenStudy (mallorysipp234):

So is the phase shift -pi/2 @mathmale

OpenStudy (mallorysipp234):

@mathmale ^^^^

OpenStudy (mathmale):

that's right! you draw the basic sine function. Then, whn you've finished, you shift the WHOLE GRAPH pI/2 units to the left. Then, as before, you shift the whole graph downward by 5 units. then you're done!

OpenStudy (mathmale):

Which answer do you believe is correct, based upon our discussion?

OpenStudy (mallorysipp234):

@mathmale would it be A?

OpenStudy (mathmale):

sorry, no, primarily because the amplitude is positive 4, not negative 4. Your a is neg, but the ampl. is |a| = |-4|=?

OpenStudy (mathmale):

enjoyed working with you!! but I need to get off OpenStudy and the 'Net and get some work done. See you again!

OpenStudy (mallorysipp234):

@mathmale so it's C, then. thank you for your help. (:

OpenStudy (anonymous):

Ty so much! This was correct (just took it)